Divide a number in such a ratio 1/2,1/3,1/9 so that total number come 18
Answer by stanbon(75887)

Divide a number in such a ratio 1/2,1/3,1/9 so that total number come 18
----
Let the numbers be x/2, x/3, and x/9
----
Equation:
x/2 + x/3 + x/9 = 18
--------
Multiply thru by 18 to get:
9x + 6x + 2x = 18^2
-----
17x = 324
x = 19.0588
------
x/2 = 9.5294
x/3 = 6.3529
x/9 = 2.1176
================
Cheers,
Stan H.
================